[figure: 9.7 cm]

DIFFUSIVE REGIME FOR CRITICAL SOUND ATTENUATION

AND DISPERSION IN MAGNETS

Andrzej Pawlak and Maciej Błaszak Institute of Physics, A. Mickiewicz University, Poznań, Poland

We consider a new regime for sound propagation in magnetic systems above the critical point. On the basis of a general statistical model including a flow of energy between spin and lattice subsystems we found that there is a region of parameters where the sound attenuation and dispersion can be expressed in a similar way as for noncritical systems for which only the heat conduction matters. Contrary to the classical heat diffusion case in the critical magnetic system both the specific-heat, entering the reduced frequency, as well as the effective amplitude are frequency and temperature dependent reflecting the critical singularities present in the system near its critical point.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Andrzej Pawlak

Address for correspondence : Umultowska 85, 61-614 Poznań, Poland

Email address : pawlak@main.amu.edu.pl